# Compilation for Dynamically Field-Programmable Qubit Arrays with Efficient and Provably Near-Optimal Scheduling

### v1

@@ VII Related Work
We have introduced the three existing works that can be directly compared with Enola: OLSQ-DPQA [25, 10], Q-Pilot [26], and Atomique [27]. In this section, we discuss other related works.
@@
A key hardware assumption of ours is that the Rydberg laser globally excites all qubits. An individually addressed Rydberg laser has also been demonstrated, but the two-qubit gate fidelity so far at 92.5% [37] is much lower than the global approach. With individual addressability, the qubits can be routed logically with SWAP gates like quantum computing platforms with a fixed coupling graph such as superconducting circuits. Baker et al. [38] covered the layout synthesis under such a hardware setting. Li et al. [39] further considered the detailed durations for different gates in the scheduling. Patel et al. [40] proposed a method of logic resynthesis to leverage three-qubit gates available on neutral atoms. The SWAPs for routing qubits can sometimes become ‘free lunch’ after the resynthesis.
Some works also utilizes the movement capabilities on neutral atoms. Brandhofer et al. [41] targeted an architecture with a more restricted kind of movement, ‘1D displacement’. Nottingham et al. [42] and Schmid et al. [43] proposed to combine the SWAP and AOD movement capabilities for routing qubits. However, these works still rely on individual addressability.
VIII Conclusion and Future Direction
@@
In this paper, we formulated three tasks in the compilation for DPQA: scheduling, placement, and routing. We presented efficient solutions to all of them and integrated these solutions in our compiler, Enola. Most notably, because the scheduling is based on the Misra&Gries edge coloring algorithm, Enola generates provably near-optimal number of two-qubit gate stages. Our placement based on simulated annealing and routing based on independent set also proves effective. This paper leads to a few promising directions.
@@
1) Synergy between the three tasks. For example, the edge coloring only assigns two-qubit gates to stages without explicitly ordering these stages. Exploring this ordering resembles placement-driven scheduling [44] and may further decrease the total movement time. Further improvement is also possible via routability-based placement [45] because the current placement only reduces the Euclidean distance of two-qubit gates without considering whether the moves corresponding to these gates are compatible.
@@
2) More detailed formulation of the routing problem. For example, the current notion of compatibility considers the entirety of movements. However, a move may not entirely be compatible with another move but becomes compatible after it progresses past a certain portion. Exploring these opportunities may also require innovation in the lower level instruction set. Separately, the NP-hardness of routing justifies further exploration in solver-based methods. Although these methods cannot solve large-scale problems, it is still valuable to pursue optimal solutions for critical and frequent subroutines or a coarsened solution in a multilevel flow.
@@
3) Application-specific compilation. This work achieves the highest improvement over previous state of the art on commutation groups of two-qubit gates. However, there are commutation relations on higher level structures like Pauli string unitaries in quantum simulation applications [46] that lies out of the scope of this paper. Another example is the decoding circuits for quantum error correction codes [47]. Although they are not commutation groups directly, there is flexibility in the ordering of two-qubit gates.
@@
4) Adaptation to hardware capabilities and co-design. DPQA with separate storage and entanglement zones have been demonstrated [9] where the Rydberg laser only illuminates the entanglement zone. This work is still useful to handle what happens inside the entanglement zone, but a higher level framework of shuttling qubits between the zones should be developed. An efficient and near-optimal compilation such as Enola also aids in the co-design of the hardware, e.g., exploring the impact of more AODs in DPQA.
@@ Acknowledgment
This work is funded by NSF grant 442511-CJ-22291. The authors would like to thank Dolev Bluvstein and Harry Hengyun Zhou for valuable discussions.
Appendix: NP-Hardness Result of the DPQA Routing Problem
In the DPQA routing problem, we are given a set of pairwise disjoint two-qubit gates and the qubit placement, i.e., a map from qubits to interaction sites. The goal is to generate compatible sets of moves to realize all the two-qubit gates. By our assumption, each move concerns one and only one gate, so the number of atom transfers is fixed regardless the routing strategy. Thus, different routing strategies have the same two-qubit gate fidelity and atom transfer fidelity, but may have different decoherence fidelity due to different time spent. An optimal routing solution should minimize both the time of AOD movements and that of atom transfers. In the main text, we focus on the former. To minimize the atom transfer time, we need to consider which qubits can be picked up in parallel. To discuss this problem, we need to introduce some terminology first.
For a binary matrix $M\in\mathbb{B}^{m\times n}$, a combinatorial rectangle is the elements $M_{ij}$ where $i\in X\subseteq[m]$ and $j\in Y\subseteq[n]$, i.e., elements at the intersections of some rows and columns. If all the elements in a rectangle are 1’s, we call it a 1-monochromatic rectangle or a rank-1 submatrix. An example is provided in Fig. 10a where different markers indicate different 1-monochromatic rectangles. The rectangular partitioning problem asks for the minimum number of 1-monochromatic rectangles to partition all the 1’s in the matrix. This problem being NP-hard is because its decision version is equivalent to the normal set basis problem which is NP-complete [29, 48].
Because of the product structure of the AOD, it can transfer qubits locating at a rectangle of SLM traps, e.g., in Fig. 1a, it aligns with 2-by-2 traps and picks up 3 qubits. If more qubits need to be collected, we can slightly shift the existing AOD rows and columns, and ramp up some other rows and columns which, again, will pick up qubits in a rectangle of SLM traps. For example, in Fig. 10a, 1’s like $M_{00}$ means the qubit is to be collected, whereas 0’s like $M_{13}$ means the qubit should not be collected. The 5 rectangles to partition the matrix incur 5 parallel pick-ups. Since each parallel pick-up takes a constant time $T_{\text{trans}}$, we would like to minimize the number of rectangles. Below, we prove that the rectangular partitioning problem is reducible to the DPQA routing problem. The idea is that in the reduced routing problem, there is only one compatible set of moves, so an optimal routing solution collects qubits with a minimum number of parallel pick-ups, i.e., rectangles.
Figure 10: NP-hardness of routing. a) A rectangular partitioning problem. b) A DPQA routing problem reduced from a).
Theorem 2.
The DPQA routing problem is NP-hard.
Proof. Given a matrix $M\in\mathbb{B}^{m\times n}$, e.g., Fig. 10a, we construct the set of two-qubit gates to route: if and only if $M_{ij}=1$, there is a gate between $q_{2(ni+j)}$ and $q_{2(ni+j)+1}$. The qubit placement is from left to right and from top to bottom. Each row contains $2n$ qubits as illustrated in Fig. 10b. The black dots are the qubits with even indices whereas the gray dots are the qubits with odd indices. Given this qubit placement, the movements for all the gates can be performed simultaneously by shifting all the black dots at the beginning of all the arrows to the right by one unit. Thus, the optimal routing solution provides the minimum number of parallel pick-ups to collect qubits corresponding to the 1’s in $M$, which is the rectangular partitioning number of $M$. Thus, we reduce an NP-hard problem, rectangular partitioning, to a DPQA routing problem, which means the DPQA routing problem is NP-hard. ∎

### v2

@@ 7. Related Work
We have introduced the three existing works that can be directly compared with Enola: OLSQ-DPQA (Tan et al., 2022, 2024a), Q-Pilot (Wang et al., 2024b), and Atomique (Wang et al., 2024a). In this section, we discuss other related works.
A key hardware assumption in our work is that the Rydberg laser globally excites all qubits. Although individually addressed Rydberg lasers have been demonstrated (Graham et al., 2022), the scale and operation fidelity of such platforms are significantly lower than that of the global approach (Bluvstein et al., 2022). With individual addressability, qubits that are idling during a Rydberg exposure will not experience the same noise level as those involved in two-qubit gates. Furthermore, moving atoms may be unnecessary when individual addressability is available, as qubits do not need to avoid each other if they are not participating in two-qubit gates during a Rydberg exposure. In this scenario, the optimization objective would shift from minimizing the number of Rydberg stages to minimizing the number of gates.
Consequently, the qubits can be routed logically using SWAP gates, similar to the methods employed in quantum computing platforms with a fixed coupling graph, such as superconducting circuits. Due to the fundamental differences in optimization objectives, most of the mentioned previous approaches (Maslov et al., 2008; Fan et al., 2022; Li et al., 2019; Wille et al., 2019; Tan and Cong, 2021a; Huang et al., 2022; Li et al., 2023a; Park et al., 2022; Zhang et al., 2021; Molavi et al., 2022; Tan and Cong, 2021b, 2020; Lin et al., 2023; Wu et al., 2022; Shaik and van de Pol, 2023) does not directly apply to the DPQA compilation problem considered in this work.
@@
Baker et al. (Baker et al., 2021) investigated layout synthesis for a fixed atom array with individually addressed Rydberg lasers. Li et al. (Li et al., 2023b) further considered the detailed durations for different gates in the scheduling process under this hardware setting. Patel et al. (Patel et al., 2022) proposed a method for logic resynthesis to leverage three-qubit gates available on neutral atoms, with the SWAPs for routing qubits sometimes becoming a ‘free lunch’ after the resynthesis.
Several works utilized movement capabilities even when assuming an individually addressed Rydberg laser. Brandhofer et al. (Brandhofer et al., 2021) targeted an architecture with a more restricted type of movement known as ‘1D displacement.’ Schmid et al. (Schmid et al., 2023) proposed a combination of SWAP and AOD movements for routing qubits. Again, due to the differing hardware assumptions, these previous approaches cannot be directly compared with the work presented here.
8. Conclusion and Future Direction
@@
In this paper, we formulated three tasks in the layout synthesis for DPQA: scheduling, placement, and routing. We presented efficient solutions to all of them and integrated these solutions in our compiler, Enola. Most notably, because the scheduling is based on the Misra-Gries edge coloring algorithm, Enola generates provably near-optimal number of two-qubit gate stages. Our placement based on simulated annealing and routing based on independent set also proves effective. This paper leads to a few promising directions.
@@
1) Synergy between the three tasks. In this work, our compilation is conducted in a single pass that encompasses scheduling, placement, and routing. However, it may be advantageous to implement further multi-pass optimizations or iterative refinement of solutions for these problems. For example, currently the edge coloring assigns two-qubit gates to stages but the order of these stages is random. Exploring this ordering resembles placement-driven scheduling (Guo et al., 2021) and may further decrease the total movement time. Further improvement is also possible via routability-based placement (Li et al., 2007) because the current placement only reduces the total Euclidean distance of two-qubit gates without considering whether the moves corresponding to these gates are compatible.
@@
2) More detailed formulation of the routing problem. For example, the current notion of compatibility considers the entirety of movements. However, a move may not entirely be compatible with another move but becomes compatible after it progresses past a certain portion. Exploring these opportunities may also require innovation in the lower-level instruction set. Separately, the NP-hardness of routing justifies further exploration in solver-based methods. Although these methods cannot solve large-scale problems, it is still valuable to pursue optimal solutions for critical and frequent subroutines or a coarsened solution in a multilevel flow.
@@
3) Application-specific compilation. This work achieves the highest improvement over previous state of the art on commutation groups of two-qubit gates. However, there are commutation relations on higher level structures like Pauli string unitaries in quantum simulation applications (Li et al., 2022) that lies out of the scope of this paper. Another example is the syndrome extraction for quantum error correcting codes, especially good quantum low-density parity-check codes (Breuckmann and Eberhardt, 2021; Xu et al., 2024). Although these circuits are not commutation groups, there is flexibility in the ordering of two-qubit gates.
@@
4) Adaptation to hardware capabilities and co-design. For example, it is possible to trap atoms in multiple 2D planes, making the architecture 3D (Schlosser et al., 2023). Another example is DPQA with separate storage and entanglement zones (Bluvstein et al., 2024) where the Rydberg laser only illuminates the entanglement zone. This work is still useful to handle what happens inside the entanglement zone, but a higher-level framework of shuttling qubits between the zones should be developed. An efficient and effective compiler such as Enola also aids in the hardware co-design by providing improvement estimations.
@@ Acknowledgment
This work is funded by NSF grant 2313083. The authors would like to thank Dolev Bluvstein, Harry Hengyun Zhou and Prof. Mikhail D. Lukin for valuable discussions.